Please help me solve this problem:
cot-1[tan(2x)] + cot-1[tan(3x)] = x

Answer by Edwin McCravy(20060)

cot-1[tan(2x)] + cot-1[tan(3x)] = x

Since tangent and cotangent have period 180°, 

cot-1[tan(2x±180°k)] + cot-1[tan(3x±180°)] = x

replace tan(2x±180°k) by cot[90°-(2x±180°k)] and tan(3x±180°k) by 
cot[90°-(3x+180°k)]

cot-1{cot[90°-(2x±180°)]} + cot-1{cot[90°-(3x±180°)]} = x

[90°-(2x±180°k)] + [90°-(3x±180°k)] = x

90° - 2x ∓ 180°k + 90° - 3x ∓ 180°k = x

                  180° ∓ 360°k - 5x = x

                       180° ∓ 360°k = 6x
                
                       180°(1 ∓ 2k) = 6x

                        30°(1 ∓ 2k) = x

Since k can take on negative integer values, we
can just use a + sign.

                        30°(2k + 1) = x

So x can be any odd multiple of 30°.

However we must discard this because the original equation contains

tan(3x), which would be undefined, for 3x would be odd multiples of 90°
and tangents of odd multiples of 90° is undefined.

So after all that, we find that there is no solution.

Edwin